# A new approach to principal–agent problems with volatility control

### v1

@@ 3.2 Solving the ‘first-best’ problem
We demonstrate in this section that the alternative ‘first-best’ problem, namely 3, can be ‘solved’ relying only on the theory of BSDEs. More precisely, one can introduce a relevant form of contracts, which will be proved to be without loss of generality, in order to reformulate the principal’s problem into a more standard stochastic control problem, defined as 4 below. In other words, while 1 is proved to be equivalent to 2 in [9] using the theory of second-order BSDEs, we prove in this section that 3 is equivalent to 4, relying on BSDEs only.
@@
In our reformulated problem, because the quadratic variation is originally fixed by the principal, the relevant form of contract will actually be similar to the one considered in principal–agent problems with drift control only. More precisely, we are led to consider here $\xi=Y_{T}^{y_{0},Z}$, where the process $Y^{y_{0},Z}$ is defined as the solution to the following ODE with random coefficients,
$Y_{t}^{y_{0},Z}=y_{0}-\int_{0}^{t}{\cal H}_{\rm A}^{\circ}\big{(}s,X,Y^{y_{0},Z}_{s},Z_{s},\Sigma_{s}\big{)}\mathrm{d}s+\int_{0}^{t}Z_{s}\cdot\mathrm{d}X_{s},\quad\mathbb{P}\textnormal{--a.s.},\;\text{for all }\mathbb{P}\in{\cal P}^{\circ}(\Sigma).$ (3.12)
for some process $Z$ satisfying appropriate integrability conditions. More precisely, in parallel with Definition 2.4, we define the set ${\cal V}^{\circ}(\Sigma)$ of appropriate processes $Z$ as follows.
@@ 2. $(ii)$
Before stating the main result of this section, we first solve the agent’s problem (3.6), in order to confirm the intuition on the relevant contract form (3.12), already supported by the previous remark. The proof of the following result only relies on the theory of BSDEs, as opposed to 2BSDEs, needed to characterise the agent’s problem originally defined in (2.5) (see [9, Proposition 4.6]).
@@
Let $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$. There exist a unique ${\cal F}_{0}^{\widebar{\mathbb{P}}+}$-measurable random variable $Y_{0}$, satisfying $\mathbb{E}^{\widebar{\mathbb{P}}}\left[Y_{0}\right]\geq R_{\rm A}$, and a unique $Z\in\mathbb{H}^{p}$ such that $\xi=Y_{T}$, for $Y\in\mathbb{D}^{p}$ defined as the solution to the following ODE:
@@
    $\mathbb{E}^{\widebar{\mathbb{P}}}\Bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}\mathrm{d}t\bigg{)}^{p^{\prime}}\Bigg{]}$ $\leq\mathbb{E}^{\widebar{\mathbb{P}}}\bigg{[}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}=\mathbb{E}^{\widebar{\mathbb{P}}}\bigg{[}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}$
    $=\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}M_{T}^{\nu}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}$
    $\leq\mathbb{E}^{\mathbb{P}^{\nu}}\Bigg{[}\bigg{(}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{)}^{q_{1}}\Bigg{]}^{\frac{1}{q_{1}}}\mathbb{E}^{\mathbb{P}^{\nu}}\Big{[}\big{|}M_{T}^{\nu}\big{|}^{q_{2}}\Big{]}^{\frac{1}{q_{2}}},$
@@
    $\mathbb{E}^{\widebar{\mathbb{P}}}\Bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}\mathrm{d}t\bigg{)}^{p^{\prime}}\Bigg{]}$ $\leq C\;\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p_{1}}\mathrm{d}t\bigg{]}^{\frac{1}{q_{1}}}\leq C\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}\big{|}c(t,X,\nu_{t})\big{|}^{p_{1}}\mathrm{d}t\bigg{]}^{\frac{1}{q_{1}}},$
@@
${\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(T)Y_{T}-\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)c(t,X,\nu_{t})\mathrm{d}t=Y_{0}$ $-\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)\big{(}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t})\big{)}\mathrm{d}t$
@@
since $M^{\nu}$ is both an $(\mathbb{F}^{\widebar{\mathbb{P}}},\widebar{\mathbb{P}})$- and an $(\mathbb{F}^{\widebar{\mathbb{P}}+},\widebar{\mathbb{P}})$-martingale (see Neufeld and Nutz (2014) [25, Proposition 2.2]). We thus have:
@@
and, since $\xi$ is admissible, the participation constraint must be satisfied, which implies $\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]\geq R_{\rm A}$. This already concludes the proof of $(i)$. Finally, recall that we implicitly assume in 3 that for any $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$, there should exist an optimal response for the agent, namely $\mathbb{P}^{\circ}\in{\cal P}^{\circ}(\Sigma)$ such that $V^{\circ}_{\rm A}(\xi,\Sigma)=J_{\rm A}(\xi,\mathbb{P}^{\circ})$. Therefore, looking back at (3.16), we deduce from the existence of $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(\xi,\Sigma)$ that
@@
with, for $\Sigma\in{\cal S}$, $y_{0}\geq R_{\rm A}$ and $Z\in{\cal V}^{\circ}(\Sigma)$,
@@ 3. $(iii)$
While it is expected that $V_{\rm P}^{\circ}\geq\widetilde{V}^{\circ}_{\rm P}$, the previous result highlights that we actually have equality. In other words, to solve the principal’s problem in the ‘first-best’ reformulation stated in 3, one can restrict the study to contracts of the form (3.12), without loss of generality. Once again, we wish to insist on the fact that the proof of the previous theorem, detailed below, mostly follows the arguments already developed in the proof of Proposition 3.8, and therefore does not rely on the theory of 2BSDEs.
@@
    Step 1. We first argue that the inequality $V_{\rm P}^{\circ}\geq\widetilde{V}^{\circ}_{\rm P}$ is straightforward, since the principal’s problem corresponding to the value $\widetilde{V}^{\circ}_{\rm P}$ is by definition restricted to contracts of the form (3.12). Indeed, as already mentioned, we clearly have that, for fixed $\Sigma\in{\cal S}$, any contract defined by $\xi=Y_{T}^{y_{0},Z}$ through (3.12) for some $y_{0}\in\mathbb{R}$ and $Z\in{\cal V}^{\circ}(\Sigma)$ is an ${\cal F}_{T}$-measurable random variable, satisfying in addition the integrability condition (${\rm I}^{p}_{\rm A}$) by definition of $Z\in{\cal V}^{\circ}$. To conclude that $\xi\in\Xi^{\circ}(\Sigma)$, it suffices to show that this contract $\xi$ satisfies in addition the agent’s participation constraint, namely $(iii)$ in the theorem. To prove this, we can follow the same arguments as in Step 2 in the proof of Proposition 3.8. More precisely, through similar computations, we obtain
@@
which directly implies that (3.13) holds for $\mathbb{P}^{\circ}$, or equivalently that $Z$ satisfies (3.13) in Definition 3.6. To summarise, for a contract $Y_{T}^{y_{0},Z}\in\Xi^{\circ}(\Sigma)$, we have $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}=y_{0}$, and this value is achieved for the optimal effort $\nu^{\mathbb{P}^{\circ}}$, defined as a maximiser of the Hamiltonian ${\cal H}_{\rm A}^{\circ}$. We conclude that both contracts $Y_{T}$ and $Y_{T}^{y_{0},Z}$ generate the same efforts, and induce the same value for the agent. Therefore, without loss of generality, the principal’s optimisation problem can be restricted to contracts of the form (3.12), namely $\xi:=Y_{T}^{y_{0},Z}$, where the process $Y^{y_{0},Z}$ is defined as the solution to ODE (3.12), for $y_{0}\geq R_{\rm A}$. It remains to verify that $Z\in{\cal V}^{\circ}(\Sigma)$, to conclude that $\widetilde{V}_{\rm P}^{\circ}=V_{\rm P}^{\circ}$. Note that $Z\in\mathbb{H}^{p}$ and satisfies Definition 3.6 $(ii)$, as mentioned above. Finally, following similar arguments as in Step 1 in the proof of Proposition 3.8, one can prove that $Y^{y_{0},Z}\in\mathbb{D}^{p}$, so that Definition 3.6 $(i)$ is satisfied. We conclude that $Z\in{\cal V}^{\circ}(\Sigma)$. ∎
@@
Finally, one can notice that, for fixed $y_{0}\geq R_{\rm A}$, $\underline{V}_{\rm P}^{\circ}(y_{0})$ corresponds to the value of a (more) standard stochastic control problem, with two state variables $X$ and $Y^{y_{0},Z}$, controlled through $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$. As done to introduce 2, one can compute the dynamics of these state variables under $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(Y^{y_{0},Z},\Sigma)$, and conclude that solving 3 is equivalent to solving 4, introduced below.
@@
$\mathrm{d}Y_{t}$ $=\Big{(}c\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}+Y_{t}k\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\mathrm{d}W_{t},\quad t\in[0,T],$ (3.20b)
@@
One can already notice that 4 stated above is extremely similar to 2. Indeed, in both problems, the principal optimises the same criteria, and, apart from the agent’s optimal effort, the dynamics of the state variables in both problems are identical. The only yet main difference between the two problems is that, in 4, this optimal effort depends (among other variables) on $\Sigma$, while, in 2, it is instead a function of $\Gamma$. We will nevertheless prove, in the next section, that there is a one-to-one correspondence between $\Sigma$ and $\Gamma$ so that optimising one or the other actually lead to the same value for the principal, i.e. $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$.
@@ 3.3 Penalisation contracts
The goal of this section is to prove the equivalence between the original problem and its ‘first-best’ reformulation, introduced by 3. As mentioned in the introduction, principal–agent problems with moral hazard are often compared to their ‘first-best’ counterpart, in which the principal can directly impose the agent’s controls. It is usually straightforward to prove that the principal’s value in the first-best scenario is greater than her value in the moral hazard case, since in the first-best case, the principal can directly choose the agent’s control. These two problems are equivalent when the principal actually observes the agent’s controls and can design a contract which strongly penalise him if he deviates from the recommended efforts.
We wish to follow the same type of reasoning here. Notice that we already proved in Lemma 3.4 that the principal’s value $V_{\rm P}^{\circ}$ in the alternative ‘first-best’ problem (3) is greater than the value $V_{\rm P}$ in the original problem (1). Therefore, it remains to design appropriate penalisation contracts to show that, starting from 1, one can actually achieve the value $V_{\rm P}^{\circ}$. This will imply the converse inequality, namely $V_{\rm P}\geq V_{\rm P}^{\circ}$, and thus conclude the proof of the main result of this paper, stated below.
@@
Solving 3 is equivalent to solving 1, i.e. $V_{\rm P}^{\circ}=V_{\rm P}$.
The idea is to start from the original problem (2.9) and restrict the study to contracts of the form (2.11) introduced in [9], hence considering 2. Then, we only have to prove that this problem is equivalent to 4. Indeed, we already have by Lemma 2.5 that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, and by Theorem 3.9 that $V^{\circ}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$. Therefore, it only remains to prove that $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$ to achieve the desired result, namely $V_{\rm P}\geq\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}=V^{\circ}_{\rm P}$.
@@
For any fixed $y_{0}\geq R_{\rm A}$, there exists a one-to-one correspondence between $\Sigma\in{\cal S}$, $Z\in{\cal V}^{\circ}(\Sigma)$ in 4 and $(Z,\Gamma)\in{\cal V}$ in 2, so that $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$. More precisely, for any $(Z,\Gamma)\in{\cal V}$ and the corresponding state variables $(X,Y)$ with dynamics given by (2.17), the corresponding process $\Sigma\in{\cal S}$ is defined by
$\Sigma_{t}\in\operatorname*{arg\,max}_{S\in\mathbb{S}_{t}(X)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},S)+\dfrac{1}{2}{\rm Tr}\big{[}\Gamma_{t}S\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$
Conversely, for any $\Sigma\in{\cal S},Z\in{\cal V}^{\circ}(\Sigma)$ and the corresponding state variables $(X,Y)$ with dynamics given by (3.20), the corresponding process $\Gamma$ is defined by
$\Gamma_{t}\in\operatorname*{arg\,min}_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,X,Y_{t},Z_{t},\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma\Sigma_{t}\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{-a.e. on }[0,T]\times\Omega.$

### v2

@@ 3.2 Solving the ‘first-best’ problem
While Problem 1 is proved to be equivalent to Problem 2 in [9] using the theory of 2BSDEs, we demonstrate in this section that the alternative ‘first-best’ problem, namely Problem 3 introduced above, can also be reformulated into a more standard stochastic control problem, defined as Problem 4 below, but relying on BSDEs only.
@@
In other words, $2{\cal H}_{\rm A}(\cdot,\gamma)$ is the convex conjugate, or Legendre–Fenchel transformation, of $-2{\cal H}^{\circ}_{\rm A}(\cdot,S)$.
In our reformulated problem, because the quadratic variation is originally fixed by the principal, the relevant form of contract will actually be similar to the one considered in principal–agent problems with drift control only. More precisely, we are led to consider here $\xi=Y_{T}^{y_{0},Z}$, where the process $Y^{y_{0},Z}$ is defined as the solution to the following SDE,
$Y_{t}^{y_{0},Z}=y_{0}-\int_{0}^{t}{\cal H}_{\rm A}^{\circ}\big{(}s,X,Y^{y_{0},Z}_{s},Z_{s},\Sigma_{s}\big{)}\mathrm{d}s+\int_{0}^{t}Z_{s}\cdot\mathrm{d}X_{s},\quad\mathbb{P}\textnormal{--a.s.},\;\text{for all }\mathbb{P}\in{\cal P}^{\circ}(\Sigma),$ (3.13)
for a constant $y_{0}\in\mathbb{R}$ and a process $Z$ satisfying appropriate integrability conditions. In parallel with Definition 2.4, we define the set ${\cal V}^{\circ}(\Sigma)$ of appropriate processes $Z$ as follows.
@@ 2. $(ii)$
Before stating the main result of this section, we first solve the agent’s problem (3.6), in order to confirm the intuition on the relevant contract form (3.13) highlighted by the previous remark. The proof of this result only relies on the theory of BSDEs, as opposed to 2BSDEs, needed to characterise the agent’s problem originally defined in (2.5) (see [9, Proposition 4.6]).
@@
Let $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$. There exist a unique ${\cal F}_{0}^{\widebar{\mathbb{P}}+}$-measurable random variable $Y_{0}$, satisfying $\mathbb{E}^{\widebar{\mathbb{P}}}\left[Y_{0}\right]\geq R_{\rm A}$, and a unique $Z\in\mathbb{H}^{p}$ such that $\xi=Y_{T}$, for $Y\in\mathbb{D}^{p}$ defined as the solution to the following SDE:
@@
    $\mathbb{E}^{\widebar{\mathbb{P}}}\Bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}\mathrm{d}t\bigg{)}^{p^{\prime}}\Bigg{]}$ $\leq\mathbb{E}^{\widebar{\mathbb{P}}}\bigg{[}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}=\mathbb{E}^{\widebar{\mathbb{P}}}\bigg{[}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c_{\rm A}(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}$
    $=\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}M_{T}^{\nu}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c_{\rm A}(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}$
    $\leq\mathbb{E}^{\mathbb{P}^{\nu}}\Bigg{[}\bigg{(}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c_{\rm A}(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{)}^{q_{1}}\Bigg{]}^{\frac{1}{q_{1}}}\mathbb{E}^{\mathbb{P}^{\nu}}\Big{[}\big{|}M_{T}^{\nu}\big{|}^{q_{2}}\Big{]}^{\frac{1}{q_{2}}},$
@@
    $\mathbb{E}^{\widebar{\mathbb{P}}}\Bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}\mathrm{d}t\bigg{)}^{p^{\prime}}\Bigg{]}$ $\leq C\;\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c_{\rm A}(t,X,u)\big{|}^{p_{1}}\mathrm{d}t\bigg{]}^{\frac{1}{q_{1}}}\leq C\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}\big{|}c_{\rm A}(t,X,\nu_{t})\big{|}^{p_{1}}\mathrm{d}t\bigg{]}^{\frac{1}{q_{1}}},$
@@
${\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(T)Y_{T}-\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)c_{\rm A}(t,X,\nu_{t})\mathrm{d}t=Y_{0}$ $-\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)\big{(}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t})\big{)}\mathrm{d}t$
@@
since $M^{\nu}$ is both an $(\mathbb{F}^{\widebar{\mathbb{P}}},\widebar{\mathbb{P}})$- and an $(\mathbb{F}^{\widebar{\mathbb{P}}+},\widebar{\mathbb{P}})$-martingale (see [25, Proposition 2.2]), and thus
@@
and, since $\xi$ is admissible, the participation constraint must be satisfied, which implies $\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]\geq R_{\rm A}$. This already concludes the proof of $(i)$. Finally, recall that we implicitly assume in Problem 3 that for any $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$, there should exist an optimal response for the agent, namely $\mathbb{P}^{\circ}\in{\cal P}^{\circ}(\Sigma)$ such that $V^{\circ}_{\rm A}(\xi,\Sigma)=J_{\rm A}(\xi,\mathbb{P}^{\circ})$. Therefore, looking back at (3.17), we deduce from the existence of $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(\xi,\Sigma)$ that
@@
where for $y_{0}\geq R_{\rm A}$, $\Sigma\in{\cal S}$, and $Z\in{\cal V}^{\circ}(\Sigma)$,
@@ 3. $(iii)$
While it is expected that $V_{\rm P}^{\circ}\geq\widetilde{V}^{\circ}_{\rm P}$, since the principal’s problem corresponding to the value $\widetilde{V}^{\circ}_{\rm P}$ is by definition restricted to contracts of the form (3.13), the previous result highlights that we actually have equality. In other words, to solve the principal’s problem in the ‘first-best’ reformulation stated in Problem 3, one can restrict the study to contracts of the form (3.13), without loss of generality. Once again, we wish to insist on the fact that the proof of the previous theorem, detailed below, mostly follows the arguments already developed in the proof of Proposition 3.8, and therefore does not rely on the theory of 2BSDEs.
@@
    Step 1. We first argue that the inequality $V_{\rm P}^{\circ}\geq\widetilde{V}^{\circ}_{\rm P}$ holds. As already mentioned, we clearly have that, for fixed $\Sigma\in{\cal S}$, any contract defined by $\xi=Y_{T}^{y_{0},Z}$ through (3.13) for some $y_{0}\in\mathbb{R}$ and $Z\in{\cal V}^{\circ}(\Sigma)$ is an ${\cal F}_{T}$-measurable random variable, satisfying in addition the integrability condition (${\rm I}^{p}_{\rm A}$) by definition of $Z\in{\cal V}^{\circ}$. To conclude that $\xi\in\Xi^{\circ}(\Sigma)$, it suffices to show that this contract $\xi$ satisfies in addition the agent’s participation constraint, namely $(iii)$ in the theorem. To prove this, we can follow the same arguments as in Step 2 in the proof of Proposition 3.8. More precisely, through similar computations, we obtain
@@
which directly implies that (3.14) holds for $\mathbb{P}^{\circ}$, or equivalently that $Z$ satisfies (3.14) in Definition 3.6. To summarise, for a contract $Y_{T}^{y_{0},Z}\in\Xi^{\circ}(\Sigma)$, we have $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}=y_{0}$, and this value is achieved for the optimal effort $\nu^{\mathbb{P}^{\circ}}$, defined as a maximiser of the Hamiltonian ${\cal H}_{\rm A}^{\circ}$. We conclude that both contracts $Y_{T}$ and $Y_{T}^{y_{0},Z}$ generate the same efforts, and induce the same value for the agent. Therefore, without loss of generality, the principal’s optimisation problem can be restricted to contracts of the form (3.13), namely $\xi:=Y_{T}^{y_{0},Z}$, where the process $Y^{y_{0},Z}$ is defined as the solution to ODE (3.13), for $y_{0}\geq R_{\rm A}$. It remains to verify that $Z\in{\cal V}^{\circ}(\Sigma)$, to conclude that $\widetilde{V}_{\rm P}^{\circ}=V_{\rm P}^{\circ}$. Note that $Z\in\mathbb{H}^{p}$ and satisfies Definition 3.6 $(ii)$, as mentioned above. Finally, following similar arguments as in Step 1 in the proof of Proposition 3.8, one can prove that $Y^{y_{0},Z}\in\mathbb{D}^{p}$, so that Definition 3.6 $(i)$ is also satisfied. ∎
@@
Finally, one can notice that, for fixed $y_{0}\geq R_{\rm A}$, $\underline{V}_{\rm P}^{\circ}(y_{0})$ corresponds to the value of a (more) standard stochastic control problem, with two state variables $X$ and $Y^{y_{0},Z}$, controlled through $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$. One can finally compute the dynamics of these state variables under $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(Y^{y_{0},Z},\Sigma)$, and conclude that solving Problem 3 is equivalent to solving Problem 4, introduced below.
@@
$\mathrm{d}Y_{t}$ $=\Big{(}c_{\rm A}\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}+Y_{t}\,k_{\rm A}\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\mathrm{d}W_{t},\quad t\in[0,T],$ (3.21b)
@@
One can already notice that Problem 4 stated above is extremely similar to Problem 2. Indeed, in both problems, the principal optimises the same criteria, and, apart from the agent’s optimal effort, the dynamics of the state variables in both problems are identical. The only yet main difference between the two problems is that, in Problem 4, this optimal effort depends (among other variables) on $\Sigma$, while, in Problem 2, it is instead a function of $\Gamma$. We will nevertheless prove, in the next section, that there is a one-to-one correspondence between $\Sigma$ and $\Gamma$ so that optimising one or the other actually lead to the same value for the principal, i.e. $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$.
@@ 3.3 Penalisation contracts
As mentioned in introduction, principal–agent problems with moral hazard are equivalent to their first-best counterpart if the principal actually observes the agent’s efforts and can design a forcing contract, which basically forces him to implement the recommended efforts. We wish to follow the same type of reasoning here, to conclude on the equivalence between the original problem and its ‘first-best’ reformulation, introduced by Problem 3. Notice that we already proved in Lemma 3.4 that the principal’s value $V_{\rm P}^{\circ}$ in the alternative ‘first-best’ problem (Problem 3) is greater than the value $V_{\rm P}$ in the original problem (Problem 1). Therefore, it remains to introduce appropriate contracts to show that, starting from Problem 1, one can actually achieve the ‘first-best’ value $V_{\rm P}^{\circ}$. This will imply the converse inequality, namely $V_{\rm P}\geq V_{\rm P}^{\circ}$, and thus conclude the proof of the main result of this paper, stated below.
@@
Under Assumptions A and B, solving Problem 3 is equivalent to solving Problem 1, i.e. $V_{\rm P}^{\circ}=V_{\rm P}$.
The idea is to start from the original problem (2.9) and restrict the study to contracts of the form (2.11) introduced in [9], hence considering Problem 2. Then, we only have to prove that this problem is equivalent to Problem 4. Indeed, we already have by Lemma 2.5 that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, and by Theorem 3.9 that $V^{\circ}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$. Therefore, it only remains to prove that $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$ to achieve the desired result, namely $V_{\rm P}\geq\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}=V^{\circ}_{\rm P}$. The previous equality is proved in two steps, first showing that $\widetilde{V}_{\rm P}\geq\widetilde{V}^{\circ}_{\rm P}$ in Lemma 3.11, and then the converse inequality in Lemma 3.13. Indeed, while the first inequality is quite straightforward to prove, the second inequality seems to involve an additional assumption on the structure of the ‘constrained’ Hamiltonian ${\cal H}_{\rm A}^{\circ}$, and formulated in Assumption B stated below. While this assumption does not appear explicitly in [9], the proof of their main result, namely Theorem 3.6 on the optimality of the contract form (2.11), may actually fail if this assumption is not verified. We refer to Remark 3.12 below for further clarification.
@@
Let $y_{0}\geq R_{\rm A}$, $(Z,\Gamma)\in{\cal V}$, and consider the corresponding state variables $(X,Y)$ with dynamics given by (2.17) and an associated optimal effort $\nu^{\star}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})$, $t\in[0,T]$. Then there exists a process $\Sigma\in{\cal S}$ defined by
$\Sigma_{t}:=\big{[}\sigma\sigma^{\top}\big{]}\big{(}t,X,u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})\big{)},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega,$
such that $Z\in{\cal V}^{\circ}(\Sigma)$ and $\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})$, $\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega$.